# Radar Spectra-Language Model for Automotive Scene Parsing

### v1

@@ Abstract
Radar sensors are low cost, long-range, and weather-resilient. Therefore, they are widely used for driver assistance functions, and are expected to be crucial for the success of autonomous driving in the future. In many perception tasks only pre-processed radar point clouds are considered. In contrast, radar spectra are a raw form of radar measurements and contain more information than radar point clouds. However, radar spectra are rather difficult to interpret. In this work, we aim to explore the semantic information contained in spectra in the context of automated driving, thereby moving towards better interpretability of radar spectra. To this end, we create a radar spectra-language model, allowing us to query radar spectra measurements for the presence of scene elements using free text. We overcome the scarcity of radar spectra data by matching the embedding space of an existing vision-language model (VLM). Finally, we explore the benefit of the learned representation for scene parsing, and obtain improvements in free space segmentation and object detection merely by injecting the spectra embedding into a baseline model.
@@ I Introduction
To address the above, we propose to train a radar spectra-language model (RSLM) for automotive scenarios, motivated by the tremendous success of vision-language models (VLMs) like CLIP [12], or DALL-E [13]. Said model can subsequently be used to query radar measurements for contents of interest using free text, a step towards understanding the semantic content of radar spectra.
@@ IV-C Radar spectra-language model
TABLE I: Comparison of top-1 and top-100 precision scores for retrieval task for original VLM, fine-tuned VLM and RSLM on CRUW dataset.
@@ IV-D Object detection and free-space segmentation with RSLM radar embeddings
Table II summarizes object detection and segmentation performance for the models described above. The columns of the table correspond to model features: “detect backbone” signifies the use of the detection backbone (MIMO pre-encoder and FPN Radar backbone) of FFT-RadNet. “radar enc” denotes the incorporation of the radar encoder from RSLM. “RSLM weights” indicates the usage of weights from the RSLM model for the radar encoder; otherwise, it is randomly initialized. “Fine-tuned enc” signifies that the radar encoder was fine-tuned during detection training; otherwise it is frozen. Results in the table exhibit performance improvements when adding frozen radar embeddings into the model architecture. This enhancement is observed in both detection and free-space segmentation tasks, as compared to the baseline model without embeddings (“frozen enc” and “fine-tuned enc” vs. “baseline”). Fine-tuning the radar encoder does not improve object detection or segmentation performance. Furthermore, the model “from-scratch” with the same architecture as the “frozen-enc” variant, exhibits slightly higher IoU scores for free-space segmentation and similar detection performance compared to the “frozen-enc” model. In contrast, models that exclusively incorporate the radar encoder component of RSLM (“only-frozen enc”, “only fine-tuned enc”), whether frozen or trained during the last 10 epochs, do not successfully accomplish the detection task.
@@
Figure 5: Detection results of FFT-RadNet (green) and our proposed network “FFT-RadNet + RLSM encoder”(blue). The bounding box prediction of FFT-RadNet is displaced w.r.t the ground truth bounding box (red), whereas the “FFT-RadNet + RLSM encoder” predictions align well with the ground truth. Confidence score equals 0.1. Left: Bounding boxes in Cartesian coordinates displayed on radar point clouds. Right: Bounding boxes projected on image.

### v2

@@ Abstract
Radar sensors are low cost, long-range, and weather-resilient. Therefore, they are widely used for driver assistance functions, and are expected to be crucial for the success of autonomous driving in the future. In many perception tasks only pre-processed radar point clouds are considered. In contrast, radar spectra are a raw form of radar measurements and contain more information than radar point clouds. However, radar spectra are rather difficult to interpret. In this work, we aim to explore the semantic information contained in spectra in the context of automated driving, thereby moving towards better interpretability of radar spectra. To this end, we create a radar spectra-language model, allowing us to query radar spectra measurements for the presence of scene elements using free text. We overcome the scarcity of radar spectra data by matching the embedding space of an existing vision-language model. Finally, we explore the benefit of the learned representation for scene retrieval using radar spectra only, and obtain improvements in free space segmentation and object detection merely by injecting the spectra embedding into a baseline model.
@@ I Introduction
To address the above, we propose an approach which a) does not need any labeled radar spectra data and b) can be used to query radar spectra for contents of interest using free text. This is a step towards understanding the semantic content of radar spectra. We propose to train a radar spectra-language model (RSLM) for automotive scenarios, motivated by the tremendous success of vision-language models (VLMs) like CLIP [12], or DALL-E [13].
@@ IV-C Radar spectra-language model
TABLE I: Comparison of top-10 and top-100 precision scores for retrieval task for original VLM, fine-tuned VLM and RSLM on CRUW dataset.
@@ IV-D Object detection and free-space segmentation with RSLM radar embeddings
Table III summarizes object detection and segmentation performance for the models described above. The columns of the table correspond to model features: “detect backbone” signifies the use of the detection backbone (MIMO pre-encoder and FPN Radar backbone) of FFT-RadNet. “radar enc” denotes the incorporation of the radar encoder from RSLM. “RSLM weights” indicates the usage of weights from the RSLM model for the radar encoder; otherwise, it is randomly initialized. “Fine-tuned enc” signifies that the radar encoder was fine-tuned during detection training; otherwise it is frozen. Results in the table exhibit performance improvements when adding frozen radar embeddings into the model architecture. This enhancement is observed in both detection and free-space segmentation tasks, as compared to the baseline model without embeddings (“frozen enc” and “fine-tuned enc” vs. “baseline”). Fine-tuning the radar encoder does not improve object detection or segmentation performance. Furthermore, the model “from-scratch” with the same architecture as the “frozen-enc” variant, exhibits slightly higher IoU scores for free-space segmentation and similar detection performance compared to the “frozen-enc” model. In contrast, models that exclusively incorporate the radar encoder component of RSLM (“only-frozen enc”, “only fine-tuned enc”), whether frozen or trained during the last 10 epochs, do not successfully accomplish the detection task.
@@
Figure 5: Detection results of FFT-RadNet (green) and our proposed network “FFT-RadNet + RLSM encoder” (blue). The bounding box prediction of FFT-RadNet is displaced w. r. t. the ground truth (red), whereas the predictions of our model align well with the ground truth. Confidence score equals 1.0. Left: Bounding boxes in Cartesian coordinates, radar point clouds displayed for reference only. Note that the models work on spectral data. Thus objects might even be predicted at locations where no radar point clouds are visible. Right: Bounding boxes projected on image.